OpenStudy (anonymous):

Verify the identity: csc x - csc x cos^2x=sinx

OpenStudy (nikato):

factor out a cscx. what do you have now?

OpenStudy (anonymous):



OpenStudy (anonymous):

One thing that always helps is to put in terms of \(\sin\) and \(\cos\). \[ \csc x-\csc x \cos^2x = \frac1{\sin x}-\frac{\cos^2x}{\sin x} \]

OpenStudy (anonymous):

So how did you get the\[\frac{ \cos ^{2}x }{ \sin x }\]

OpenStudy (anonymous):

Oh, nvm. What would you do after that step?

OpenStudy (anonymous):

Well, it's a fraction, so you can combine them. Fortunately the denominator is already the same, so it works out easily.

OpenStudy (anonymous):

\[ \frac1{\sin x}-\frac{\cos^2x}{\sin x} = \frac{1-\cos^2x}{\sin x} \]

OpenStudy (anonymous):

At this point, you need to remember the Pythagorean identity. \[ \sin^2x+\cos^2x=1 \]

OpenStudy (anonymous):

Do you think you would be able to do the rest?

OpenStudy (anonymous):

Not really. I don't get how the pythagorean identity applies to the next step.

OpenStudy (anonymous):

\[ \frac{1-\cos^2x}{\sin x}=\frac{(\sin^2x+\cos^2x)-\cos^2x}{\sin x} \]

OpenStudy (anonymous):

That is how it will apply.

OpenStudy (anonymous):

Oh, okay. I get it. Thank you.